It lives.
Bipolar PID loop - check
Laser Inhibit - check
Status LED's - check
Current limit - to be completed.
Frank, I had trouble with the NTC op-amp. You had fitted 100R instead of 100K in the negative feedback loop. Now corrected, while it does work, the output is higher than you expected and clips at 4.30V above around 25c. Do you have a board you can check?
Also, for some reason, I have schematic V5 and the part designations do not match the last component overlay you posted. Can you please provide correct matching schematic/overlays please.
